OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x – 5y = –3 6x – 3y = –3

OpenStudy (anonymous):

@im here just to see ur question :) :) sorry don't hope from me

OpenStudy (anonymous):

\[x =(3y-3)/6\] \[2\times(3y-3)/6 - 5y = -3\] \[y-1-5y = -3 \implies -4y = -2 \implies y = 1/2\] x =1/6(3/2-3) = -1/4

OpenStudy (anonymous):

y=-3-6x/-3